[Torrey, Jeff]
The second AOM was not working, and this was resolved by unplugging the BNC on the AOM and replugging it.
There was no 775 light reaching the second filter cavity, so we investigated the second AOM on the SHG sled. After confirming that it was receiving RF power correctly by swapping cables with the first AOM, we removed the cover and visually inspected the crystal to ensure it looks good. We added a pickoff mirror and a camera after the AOM and could see two spots and fiddling with alignment for a while we could still not get the power of the frequency shifted beam to an acceptable amount (we want to see roughly 50/50). Finally after unplugging and replugging the BNC the two spots were about equal.
We then had to realign the fiber coupler. We left it at
76 uW in, 40uW out -> 52% efficiency